# Generalized saturation game

### v1

@@ 1 Introduction
One of the earliest combinatorial games on graphs is Hajnal’s triangle game in which two players alternatingly claim edges of the complete graph $K_{n}$ such that the graph $G$ of claimed edges cannot contain a triangle. The player who cannot move on his turn, loses. This innocent looking game turned out to be quite hard to analyze, so Füredi, Reimer, and Seress [8] introduced its quantitative version. The rules of the game remain the same, but the aims of the players are different. The score of the game is the total number of edges claimed by the players during the game. One of the players, Max, targets to maximize the score, while the other player, Mini, aims to minimize the score. The game saturation number is the score of the game when both players play according to an optimal strategy. Note that this might depend on who plays first, so precisely we have a Max-start and a Mini-start saturation number. Recent improvements on obtaining bounds on these numbers were obtained by Biró, Horn, and Wildstrom [2]. Clearly, the rules and the scores of this game can be changed to forbid any graph $F$, or a family of graphs $\mathcal{F}$. The game in this generality was introduced in [3] and results of this type were obtained in [5, 12, 9, 16].
@@
In this paper, we study the game variant, first proposed in [13], of the above mentioned generalized problem. The $H$-score of an $F$-saturation game is the total number of copies of $H$ at the end of the game. The $H$-score of the game when both players play optimally and Max starts is denoted by $s_{1}(n,\#H,F)$, and by $s_{2}(n,\#H,F)$ when Mini starts. When we have a conclusion that is valid independently of who starts the game, then we may simply write $s(n,\#H,F)$
@@
    If $s\geq 7$, then for any $\ell\geq 3$, we have $s(n,\#S_{\ell},P_{s})=\Theta(n^{\ell-1})$.
@@
${s(n,P_{6},S_{4})=0}.$
@@
Finally, we address $P_{5}$-saturation games with scores being the number of triangles or paths of lenght $3$.
@@ 2 Paths vs stars
In this section, we prove Theorem 1.1. First we show part (1). Observe that the number of $S_{\ell}$’s in a graph $G$ is $\sum_{v\in V(G)}\binom{\deg_{G}(v)}{\ell-1}$. Therefore, the upper bound $s(n,\#S_{\ell},P_{s})=O_{\ell,s}(n^{\ell-1})$ follows from the result of Erdős and Gallai that states $\mathop{}\!\mathrm{ex}(n,P_{s})\leq\frac{s-2}{2}n$ and so $\sum_{v\in V(G)}\deg_{G}(v)\leq(s-2)|V(G)|$ for any $P_{s}$-free graph $G$.
@@
Suppose $G_{n}$ contains a component $C$ and a vertex $v\in V(C)$ such that $C\setminus\{v\}$ has at least three components $C_{1},C_{2},C_{3}$ with $C_{i}\cup\{v\}$ containing a path of length $t$. Suppose the number of Hamiltonian components of $G_{n}$ of size at most $t$ is $T$. Then Max in the Mini-start $(G_{n},P_{j})$-game can achieve that $v$ has degree at least $T/3$ at the end of the game.
@@
Let $K^{1},K^{2},\dots,K^{m_{i}}$ be those Hamiltonian components of $G_{n}$ of size at most $t$ to which no adjacent edges have been played by either of the players before Max’s move in round $i$. We claim that in his next move, Max can join any vertex $u$ of the largest $K^{h}$ to $v$. As Mini can touch at most two components per round, if we can prove the above claim, then the statement of the lemma follows.
We prove the following statement by induction on the number of rounds passed: after Max’s move, if $C^{i}$ denotes the component containing $v$ of the game graph after round $i$, then there are at least three components $C^{\prime}$ of $C^{i}\setminus\{v\}$ such that $C^{\prime}\cup\{v\}$ contains a path ending in $v$ that has strictly more vertices than any $K^{j}$. This is certainly true at the beginning of the game as this is the assumption of the lemma. So suppose the statement is true after round $i$. Then Mini can play an edge that keeps the game graph $P_{j}$-free and after her move, there are at least two components of $C^{i}\setminus\{v\}$ that together with $v$ contain a path ending in $v$ of length longer than a largest Hamiltonian component $K^{h}$. Now Max tries to play an edge $vw$ with $w\in K^{h}$. We need to show that the game graph remains $P_{j}$-free. Suppose not. Then a copy of $P_{j}$ is created that uses $vw$ and possibly some part of a Hamiltonian path in $K^{h}$. Observe that $P_{j}$ can use at most one other component of $C^{i}\setminus\{v\}$, so the part of the copy of $P_{j}$ on the Hamiltonian path on $K^{h}$ can be replaced with the path on the non-used component ending in $v$. This is another copy of $P_{j}$ that was already present after Mini’s move. This contradiction completes the proof of the lemma. ∎
@@
Let $2\leq S\leq L<s$ be positive integers and $\varepsilon>0$ a positive real. Then in the $(E_{n},P_{s})$-game Max can either achieve that there exists a vertex in the game graph with degree $\varepsilon\frac{n}{3}$, or that the game graph has $\frac{(1-\varepsilon)n}{S+L}$ Hamiltonian components, each of sizes $S$ and $L$.
@@
Finally, if after Max finishes a stage there are no isolated edges, then Max claims an isolated edge, and Mini claims another isolated edge, as otherwise Max creates an $S_{4}$ and applies Lemma 2.1. If $M=3$, then Max extends his edge to a $P_{3}$ which Mini has to close into a triangle as otherwise Max can create an $S_{4}$, or a $P_{l}$, $l\in\{4,5\}$, with a hanging edge (depending on whether Mini touched $P_{3}$ or not), and apply Lemma 2.1. So if $M=3$, then stage $r+1$ is finished.
If $M\geq 4$, then Max connects his first edge to $e$ to obtain a $P_{4}$ that Mini next has to close to a $C_{4}$ as otherwise Max can connect an isolated vertex to the neighbor of the degree one vertex of the $P_{4}$ and apply Lemma 2.1. From here on, Max can follow the same strategy as above to obtain his Hamiltonian component of size $M$ and finish stage $r+1$. ∎
@@
Note that at the end of this phase, the game graph contains at least $\frac{n}{2(S+L)}$ nice components, each of sizes $L$ and $S$, so the number of such components is at least 5 times the number of vertices that do not belong to nice components (with room to spare). We call these vertices bad. Also, in the remainder of the proof, we assume $s\geq 11$ or $s=9$ which implies $3S\geq s$. We will deal with the cases $s=7,8,10$ separately.
@@
At the beginning of Phase II, all components are Hamiltonian as nice components are such by definition, and Lemma 2.3 ensures that all other components are isolated vertices with the exception of possibly one isolated edge. During this phase, components may merge. At any point of the game, a component that contains a bad vertex is a bad component. If Max connects two nice components, then we call the resulting component a joined component. A joined component $C$ becomes a problematic component if Mini connects it to a bad component $C^{\prime}$ with $|C|+|C^{\prime}|\geq s$. A component $C$ is good if no nice component can be connected to $C$ at any of its vertices. (Observe that a component can be good and bad at the same time.)
@@
Next, suppose Mini connects the joined component $C$ to a bad Hamiltonian component $C^{\prime}$ with $|C|+|C^{\prime}|\geq s$. Let $C_{1},C_{2}$ be the nice components joined in $C$. Let $uv$ be the edge joining them with $v\in C_{1},u\in C_{2}$. As $C_{1}$ and $C_{2}$ are cliques (otherwise, as we already concluded, Max would not have been forced to create $C$) the edge $e$ connecting $C$ and $C^{\prime}$ must contain $u$ or $v$, since otherwise $e$ would create a $P_{s}$. Without loss of generality, we can assume $e=vw$ with $w\in C^{\prime}$.
@@
Now we can assume $s=4k+1$ or $s=4k+2$ and $|C^{\prime}|<S$. Then, $|C_{1}|=|C_{2}|=S$, and Max claims an edge $vu_{3}$ with $u_{3}\in C_{3}$ a nice component of size $L$ (a legal move as the largest path goes from $C_{3}$ to $C_{2}$ and contains $L+1+S<s$ vertices). Now if Mini does not claim an edge connecting $C_{2}$ and $C_{3}$, then in his next move Max claims an edge $vu_{4}$ with $u_{4}\in C_{4}$ a nice component of size $S$ and applies Lemma 2.1 with $v$ and $t=S$ ($C_{2},C_{3},C_{4}$ are components in $C\setminus\{v\}$), to obtain a vertex of degree $\Theta_{s}(n)$. Therefore, Mini claims an edge $e$ connecting $C_{2}$ and $C_{3}$. Observe that $e=uu_{3}$ as all other edges would yield a path on $L+S+S\geq s$ vertices (remember, all nice components are cliques by now).
From here on Max will always claim an edge $u_{h}x$ with $x$ being one of $u,v$ or the previous $u_{i}$’s, and $u_{h}$ a vertex of a nice component $C_{h}$ of size $S$, where $h$ is always one larger than the previously used index. The fact that this is a legal move can always be seen by applying Lemma 2.2. Note that the set $U=\{u_{1}=v,u_{2}=u,u_{3},u_{4},\dots\}$ disconnects the $C_{i}$’s, so a longest path from $C_{i}$ to $C_{j}$ contains $|C_{i}|+|C_{j}|+a_{ij}$ vertices where $a_{ij}$ is the most number of inner vertices a path from $u_{i}$ to $u_{j}$ in $G[U]$ may contain. If Mini responds by claiming an edge $e$, then $e=u_{h}y$ with $y\in U$. Indeed, connecting $C_{h}$ to a vertex in $C_{i}\setminus U$ or a vertex from $U\setminus\{x\}$ to a vertex $C_{h}\setminus\{u_{h}\}$ would create a $P_{s}$, while not connecting $C_{h}$ to any $U\setminus\{x\}$ allows Max to connect a $C_{h+1}$ to $x$ and apply Lemma 2.1 to obtain a linear degree vertex.
Let us take a closer look at how Max will connect new components. Max first claims the edge $u_{3}u_{4}$, $u_{4}\in C_{4}$, where $C_{4}$ is a nice component of size $S$. If Mini connects $u_{4}$ to $u_{1}$ or $u_{2}$, then a $P_{4k+1}$ is created from $C_{4}$ to $C_{3}$. So, if $s=4k+1$, this is not possible, and hence with his next move Max can claim $u_{3}u_{5}$ and apply Lemma 2.1 to obtain a vertex of linear degree.
If $s=4k+2$, then Mini has to claim either $u_{4}u_{1}$ or $u_{4}u_{2}$, say she claims $u_{4}u_{2}$. Then Max claims $u_{5}u_{3}$, and then Mini either does not connect $C_{5}$ to $U$ and lets Max claim $u_{3}u_{6}$ and obtain a linear degree vertex via Lemma 2.1 or her only possibility to connect $u_{5}$ to $u_{2}$ as connecting to $u_{1}$ or $u_{4}$ would create a $P_{s}$ from $C_{5}$ to $C_{3}$. But now the problematic component $C$ has become good, as connecting any further nice component to $C$ would create a $P_{s}$.
@@
Let $uv$ be the edge claimed by Mini connecting the two nice components with $v$ being in a component of size $L$. Then Max claims an edge $vw$ where $w$ is a vertex of a nice component of size $S$. This is a valid move as in the newly formed component the longest path contains at most $L+S<s$ vertices. Now Max can apply Lemma 2.1 with $v$ and $t=S$ to obtain a vertex of degree at least $\frac{1}{3}\frac{3n}{10(S+L)}=\Theta_{s}(n)$.
@@
Max can apply Lemma 2.3 with $S=L=s-1$ and $\varepsilon$ arbitrarily small to obtain pairwise vertex-disjoint Hamiltonian components of size $s-1$ that cover almost all vertices. These components cannot be joined either to each other or to other vertices as a $P_{s}$ would be created, so at the end of the game, they should all be cliques yielding the lower bound. ∎
@@
If there is no isolated vertex in the game graph, then the game is over as joining two isolated edges would create a $P_{4}$. If there is an isolated vertex, and an isolated edge, then Mini joins these, and Max can make this $P_{3}$ a triangle, but then the game is over with all components at the end are either edges or triangles. A last subcase is when there is an isolated vertex, but no isolated edge in which case, the game is over with an induced triangle matching.
@@
At the moment when Mini cannot follow her strategy from the first phase, the components of the game graph are $K_{p}$’s with $p\leq 4$ with at most one isolated vertex. Mini will join two smallest such components. After this move, because of the $P_{5}$-free property, the only components that can be merged are isolated edges and possibly a single copy of $P_{3}$. Whenever Max connects two isolated edges, Mini will make that component a $C_{4}$ thus making it unable to be connected to any other components. The only other option for Max is to connect an isolated edge $uv$ to the middle vertex $y$ of the $P_{3}$ $xyz$, say Max takes the edge $vy$. Then Mini takes the edge $uy$ and makes this component non-connectable to any other components. If later in the game, Mini has to join two isolated edges, then that $P_{4}$ is not joinable to any further components. Therefore, at the end of the game, all components will have size at most 4 and at most one component of size 5.
@@
At the moment when Mini cannot follow her strategy from the first phase, the components of the game graph are $K_{p}$’s with $p\leq 5$ with at most one isolated vertex. Mini will join the two smallest such components. As two triangles cannot be joined, if there is at most one component of size smaller than 3 before Mini’s move, then Mini joins this component to a larger one, and the final components are established, all of size at most 5. So Mini either joins an isolated vertex to an isolated edge to obtain a $P_{3}$ or two isolated edges to obtain a $P_{4}$. We will call this component the current component, and in general, there will be exactly one component, where Mini makes the first move in the second phase. Also, in this component, the players claim edges alternatingly, although not necessarily in a consecutive manner. Mini can achieve this, as Max may join two $K_{2}$’s or a $K_{2}$ and a $K_{3}$, but in this case, Mini makes this component Hamiltonian with 2 or 4 edges remaining to be claimed, and Mini only claims one such edge, if Max claims one.
@@
Subcase III/A With her first move in the second phase, Mini creates a $P_{4}$ $uvwz$
If Max then plays an edge within this component, then Mini makes this component Hamiltonian, and as it is of size 4, it is not joinable to further components. When all edges are claimed, then Mini again joins two isolated edges. If Max joins the $P_{4}$ to another component, then because of the $P_{6}$-free property, it must be an isolated edge $xy$ and the only possible connection, up to isomorphism, is the edge $xv$. Then Mini plays $xw$ and obtains a maximal $P_{6}$-free component.
Subcase III/B With her first move in the second phase, Mini creates a $P_{3}$ $uvw$
@@
If Max joins $P_{3}$ to a triangle $xyz$, then the only way to join them (up to isomorphism) is by claiming the edge $xv$. Then, Mini can claim the edge $vy$ and after that, knowing that there are no further isolated vertices, the current component cannot be joined to further components. After that Mini gets back to Subcase III/A.
If Max joins $P_{3}$ to a $K_{2}$ such that the current component becomes a $P_{5}$, then Mini closes this $P_{5}$ to a $C_{5}$, and the current component cannot be joined to further components. After claiming the remaining 5 edges of this component, Mini will be back to Subcase III/A.
Finally, suppose Max connects $P_{3}$ to an isolated edge $xy$ by claiming $xv$. Then Mini plays $yu$. It can be checked that the current component cannot be joined to any further components, so after playing the remaining 5 edges of the component, Mini is back to Subcase III/A. ∎

### v2

@@ 1 Introduction
One of the earliest combinatorial games on graphs is Hajnal’s triangle game in which two players alternatingly claim edges of the complete graph $K_{n}$ such that the graph $G$ of claimed edges cannot contain a triangle. The player who cannot move on his turn, loses. This innocent looking game turned out to be quite hard to analyze, so Füredi, Reimer, and Seress [8] introduced its quantitative version. The rules of the game remain the same, but the aims of the players are different. The score of the game is the total number of edges claimed by the players during the game. One of the players, Max, targets to maximize the score, while the other player, Mini, aims to minimize the score. The game saturation number is the score of the game when both players play according to an optimal strategy. Note that this might depend on who plays first, so precisely we have a Max-start and a Mini-start saturation number. Recent improvements on obtaining bounds on these numbers were obtained by Biró, Horn, and Wildstrom [2]. Clearly, the rules and the scores of this game can be changed to forbid any graph $F$, or a family of graphs $\mathcal{F}$. The game in this generality was introduced in [3] and results of this type were obtained in [5, 9, 12, 13, 17].
@@
In this paper, we study the game variant, first proposed in [14], of the above mentioned generalized problem. The $H$-score of an $F$-saturation game is the total number of copies of $H$ at the end of the game (each subgraph isomorphic to $H$ is connected once). The $H$-score of the game when both players play optimally and Max starts is denoted by $s_{1}(n,\#H,F)$, and by $s_{2}(n,\#H,F)$ when Mini starts. When we have a conclusion that is valid independently of who starts the game, then we may simply write $s(n,\#H,F)$.
@@
    If $s\geq 7$, then for any $\ell\geq 3$, we have $s(n,\#S_{\ell},P_{s})=\Theta_{s}(n^{\ell-1})$.
@@
${s(n,\#P_{6},S_{4})=0}.$
Here we do not discuss the result for the score of the game where we count the number of $P_{4}$’s, as this result will be discussed later in Section 3.2.
@@
Finally, we address $P_{5}$-saturation games with scores being the number of triangles or paths of length $3$.
@@ 2 Paths vs stars
In this section, we prove Theorem 1.1. First we show part (1). Observe that the number of $S_{\ell}$’s in a graph $G$ is $\sum_{v\in V(G)}\binom{\deg_{G}(v)}{\ell-1}$. Therefore, the upper bound $s(n,\#S_{\ell},P_{s})=O_{\ell,s}(n^{\ell-1})$ follows from the result of Erdős and Gallai [7] that states $\mathop{}\!\mathrm{ex}(n,P_{s})\leq\frac{s-2}{2}n$ and so $\sum_{v\in V(G)}\deg_{G}(v)\leq(s-2)|V(G)|$ for any $P_{s}$-free graph $G$.
@@
Suppose $G_{n}$ contains a component $C$ and a vertex $v\in V(C)$ such that $C\setminus\{v\}$ has at least three components $C_{1},C_{2},C_{3}$ with $C_{i}\cup\{v\}$ containing a path of length $t$ that ends in $v$. Suppose the number of Hamiltonian components of $G_{n}$ of size at most $t$ is $T$. Then Max in the Mini-start $(G_{n},P_{j})$-game can achieve that $v$ has degree at least $T/3$ at the end of the game.
@@
Let $K^{1},K^{2},\dots,K^{m_{i}}$ be those Hamiltonian components of $G_{n}$ of size at most $t$ to which no adjacent edges have been played by either of the players before Max’s move in round $i$. We claim that in his next move, Max can join any vertex $u$ of a largest $K^{h}$ to $v$. As Mini can touch at most two components per round, if we can prove the above claim, then the statement of the lemma follows.
We prove the following statement by induction on the number of rounds passed: after Max’s move, if $C^{i}$ denotes the component containing $v$ of the game graph after round $i$, then there are at least three components $C^{\prime}$ of $C^{i}\setminus\{v\}$ such that $C^{\prime}\cup\{v\}$ contains a path ending in $v$ that has strictly more vertices than any $K^{j}$. This is certainly true at the beginning of the game as this is the assumption of the lemma. So suppose the statement is true after round $i$. Then Mini can play an edge that keeps the game graph $P_{j}$-free and after her move, there are at least two components of $C^{i}\setminus\{v\}$ that together with $v$ contain a path ending in $v$ with strictly more vertices than a largest Hamiltonian component $K^{h}$.
Now Max plays an edge $vu$ with $u\in K^{h}$, and we need to show that the game graph remains $P_{j}$-free. Suppose not. Then a copy of $P_{j}$ is created that uses $vu$ and possibly some part of a Hamiltonian path in $K^{h}$. Observe that $P_{j}$ can use at most one other component of $C^{i}\setminus\{v\}$, so the part of the copy of $P_{j}$ on the Hamiltonian path on $K^{h}$ can be replaced with the path on the non-used component ending in $v$. This is another copy of $P_{j}$ that was already present after Mini’s move. Hence, Max can play as described.
That move of Max adds the third component containing a path ending in $v$, with strictly more vertices than any $K^{j}$. This concludes the inductive step, and the proof of the lemma. ∎
@@
Let $2\leq S\leq L<s$ be positive integers and $\varepsilon>0$ a positive real. Then in the $(E_{n},P_{s})$-game, where $E_{n}$ is an empty graph, Max can either achieve that there exists a vertex in the game graph with degree $\varepsilon\frac{n}{3}$, or that the game graph has $\frac{(1-\varepsilon)n}{S+L}$ Hamiltonian components, each of sizes $S$ and $L$.
@@
Finally, if after Max finishes a stage there are no isolated edges, then Max claims an isolated edge $e_{1}$, and Mini claims another isolated edge $e_{2}$, as otherwise Max creates a copy of $S_{4}$ and applies Lemma 2.1.
If $M=3$, then Max extends his edge to a copy of $P_{3}$ which Mini has to close into a triangle, as otherwise Max can create a copy of $S_{4}$, or a $P_{l}$, $l\in\{4,5\}$, with a hanging edge (depending on whether Mini touched $P_{3}$ or not), and apply Lemma 2.1. So, if $M=3$, then stage $r+1$ is finished.
If $M\geq 4$, then Max connects edges $e_{1}$ and $e_{2}$ to obtain a copy of $P_{4}$ that Mini next has to close to a $C_{4}$, as otherwise Max can connect an isolated vertex to the neighbor of a degree one vertex of the copy of $P_{4}$ and apply Lemma 2.1. From here on, Max can follow the same strategy as above to obtain his Hamiltonian component of size $M$ and finish stage $r+1$. ∎
@@
Note that at the end of this phase, the game graph either contains a vertex of high degree (in which case we are done) or at least $\frac{n}{2(S+L)}$ nice components, each of sizes $L$ and $S$, so the number of such components is at least 5 times the number of vertices that do not belong to nice components (with room to spare). We call these vertices bad. In the remainder of the proof, w.l.o.g. we assume the existence of the nice components and that $s\geq 11$ or $s=9$ which implies $3S\geq s$. We will deal with the cases $s=7,8,10$ separately.
@@
At the beginning of Phase II, all components are Hamiltonian as nice components are such by definition, and the proof of Lemma 2.3 ensures that all other components are isolated vertices with the exception of possibly one isolated edge. During this phase, components may merge. At any point of the game, a component that contains a bad vertex is a bad component. If Max connects two nice components, then we call the resulting component a joined component. A joined component $C$ becomes a problematic component if Mini connects it to a bad component $C^{\prime}$ with $|C|+|C^{\prime}|\geq s$. A component $C$ is good if no nice component can be connected to $C$ at any of its vertices. (Observe that a component can be good and bad at the same time.)
@@
Next, suppose Mini connects the joined component $C$ to a bad Hamiltonian component $C^{\prime}$ with $|C|+|C^{\prime}|\geq s$. If that happens, we still refer to this newly formed component as $C$. Similarly, later in the proof, adding new components to $C$ will cause this component to grow. Let $C_{1},C_{2}$ be the nice components joined in $C$. Let $uv$ be the edge joining them with $v\in C_{1},u\in C_{2}$. As $C_{1}$ and $C_{2}$ are cliques (otherwise, as we already concluded, Max would not have been forced to create $C$) the edge $e$ connecting $C$ and $C^{\prime}$ must contain $u$ or $v$, since otherwise $e$ would create a copy of $P_{s}$. Without loss of generality, we can assume $e=vw$ with $w\in C^{\prime}$.
@@
Now we can assume $s=4k+1$ or $s=4k+2$ and $|C^{\prime}|<S$. Then, $|C_{1}|=|C_{2}|=S$, and Max claims an edge $vu_{3}$ with $u_{3}\in C_{3}$ a nice component of size $L$ (a legal move as the largest path goes from $C_{3}$ to $C_{2}$ and contains $L+1+S<s$ vertices). Now if Mini does not claim an edge connecting $C_{2}$ and $C_{3}$, then in his next move Max claims an edge $vu_{4}$ with $u_{4}\in C_{4}$ a nice component of size $S$ and applies Lemma 2.1 with $v$ and $t=S$ ($C_{2},C_{3},C_{4}$ are all of size at least $S$), to obtain a vertex of degree $\Theta_{s}(n)$. Therefore, assume that Mini claims an edge $e$ connecting $C_{2}$ and $C_{3}$. Observe that $e=uu_{3}$ as all other edges would yield a path on $L+S+S\geq s$ vertices (remember, all nice components are cliques by now).
From here on Max will always claim an edge $u_{h}x$ with $x$ being one of $u,v$ or the previous $u_{i}$’s, and $u_{h}$ a vertex of a nice component $C_{h}$ of size $S$, where $h$ is always one larger than the previously used index. The fact that this is a legal move can always be seen by applying Lemma 2.2. Note that the set $U=\{u_{1}=v,u_{2}=u,u_{3},u_{4},\dots\}$ disconnects the $C_{i}$’s, so a longest path from $C_{i}$ to $C_{j}$ contains $|C_{i}|+|C_{j}|+a_{ij}$ vertices where $a_{ij}$ is the most number of inner vertices a path from $u_{i}$ to $u_{j}$ in $G[U]$ may contain. If Mini responds by claiming an edge $e$, then $e=u_{h}y$ with $y\in U$. Indeed, connecting $C_{h}$ to a vertex in $C_{i}\setminus U$ or a vertex from $U\setminus\{x\}$ to a vertex $C_{h}\setminus\{u_{h}\}$ would create a copy of $P_{s}$, while not connecting $C_{h}$ to any vertex of $U\setminus\{x\}$ allows Max to connect a $C_{h+1}$ to $x$ and apply Lemma 2.1 to obtain a linear degree vertex.
Let us take a closer look at how Max will connect new components. Max first claims the edge $u_{3}u_{4}$, $u_{4}\in C_{4}$, where $C_{4}$ is a nice component of size $S$. If Mini connects $u_{4}$ to $u_{1}$ or $u_{2}$, then a copy of $P_{4k+1}$ is created from $C_{4}$ to $C_{3}$. So, if $s=4k+1$, this is not possible, and hence with his next move Max can claim $u_{3}u_{5}$ and apply Lemma 2.1 to obtain a vertex of linear degree.
If $s=4k+2$, then Mini has to claim either $u_{4}u_{1}$ or $u_{4}u_{2}$, say she claims $u_{4}u_{2}$. Then Max claims $u_{5}u_{3}$, and then Mini either does not connect $C_{5}$ to $U$ and lets Max claim $u_{3}u_{6}$ and obtain a linear degree vertex via Lemma 2.1 or her only possibility is to connect $u_{5}$ to $u_{2}$ as connecting to $u_{1}$ or $u_{4}$ would create a copy of $P_{s}$ from $C_{5}$ to $C_{3}$. But now the problematic component $C$ has become good, as connecting any further nice component to $C$ would create a copy of $P_{s}$.
@@
Let $uv$ be the edge claimed by Mini connecting the two nice components with $v$ being in a component of size $L$. Then Max claims an edge $vw$ where $w$ is a vertex of a nice component of size $S$. This is a valid move as in the newly formed component the longest path contains at most $L+S<s$ vertices. Now, similarly to previous cases, Max can apply Lemma 2.1 with $v$ and $t=S$ to obtain a vertex of degree at least $\Theta_{s}(n)$.
@@
Max can apply Lemma 2.3 with $S=L=s-1$ and $\varepsilon$ arbitrarily small to obtain pairwise vertex-disjoint Hamiltonian components of size $s-1$ that cover almost all vertices. These components cannot be joined either to each other or to other vertices as a copy of $P_{s}$ would be created, so at the end of the game, they should all be cliques yielding the lower bound. ∎
@@
If there is no isolated vertex in the game graph, then the game is over as joining two isolated edges would create a copy of $P_{4}$. If there is an isolated vertex, and an isolated edge, then Mini joins these, and Max can make this copy of $P_{3}$ a triangle, but then the game is over with all components at the end are either edges or triangles. A last subcase is when there is an isolated vertex, but no isolated edge in which case, the game is over with an induced triangle matching.
@@
At the moment when Mini cannot follow her strategy from the first phase, the components of the game graph are $K_{p}$’s with $p\leq 4$ with at most one isolated vertex. Mini will join two smallest such components. After this move, because of the $P_{5}$-free property, the only components that can be merged are isolated edges and possibly a single copy of $P_{3}$. Whenever Max connects two isolated edges, Mini will make that component a copy of $C_{4}$ thus making it unable to be connected to any other components. The only other option for Max is to connect an isolated edge $uv$ to the middle vertex $y$ of the $P_{3}$ $xyz$, say Max takes the edge $vy$. Then Mini takes the edge $uy$ and makes this component non-connectable to any other components. If later in the game, Mini has to join two isolated edges, then that copy of $P_{4}$ is not joinable to any further components. Therefore, at the end of the game, all components will have size at most 4 and at most one component of size 5.
@@
At the moment when Mini cannot follow her strategy from the first phase, the components of the game graph are $K_{p}$’s with $p\leq 5$ with at most one isolated vertex. Mini will join the two smallest such components. As two triangles cannot be joined, if there is at most one component of size smaller than 3 before Mini’s move, then Mini joins this component to a larger one, and the final components are established, all of size at most 5. So Mini either joins an isolated vertex to an isolated edge to obtain a copy of $P_{3}$ or two isolated edges to obtain a copy of $P_{4}$. We will call this component the current component, and in general, there will be exactly one component, where Mini makes the first move in the second phase. Also, in this component, the players claim edges alternatingly, although not necessarily in a consecutive manner. Mini can achieve this, as Max may join two copies of $K_{2}$ or a copy of $K_{2}$ and a copy of $K_{3}$, but in this case, Mini makes this component Hamiltonian with 2 or 4 edges remaining to be claimed, and Mini only claims one such edge, if Max claims one.
@@
Subcase III/A With her first move in the second phase, Mini creates a copy of $P_{4}$ $uvwz$
If Max then plays an edge within this component, then Mini makes this component Hamiltonian, and as it is of size 4, it is not joinable to further components. When all edges are claimed, then Mini again joins two isolated edges. If Max joins the copy of $P_{4}$ to another component, then because of the $P_{6}$-free property, it must be an isolated edge $xy$ and the only possible connection, up to isomorphism, is the edge $xv$. Then Mini plays $xw$ and obtains a maximal $P_{6}$-free component.
Subcase III/B With her first move in the second phase, Mini creates a copy of $P_{3}$ $uvw$
@@
If Max joins a copy of $P_{3}$ to a triangle $xyz$, then the only way to join them (up to isomorphism) is by claiming the edge $xv$. Then, Mini can claim the edge $vy$ and after that, knowing that there are no further isolated vertices, the current component cannot be joined to further components. After that Mini gets back to Subcase III/A.
If Max joins a copy of $P_{3}$ to a $K_{2}$ such that the current component becomes a copy of $P_{5}$, then Mini closes this component to a copy of $C_{5}$, and the current component cannot be joined to further components. After claiming the remaining 5 edges of this component, Mini will be back to Subcase III/A.
Finally, suppose Max connects a copy of $P_{3}$ to an isolated edge $xy$ by claiming $xv$. Then Mini plays $yu$. It can be checked that the current component cannot be joined to any further components, so after playing the remaining 5 edges of the component, Mini is back to Subcase III/A. ∎
Now to get upper bounds in Theorem 1.1, parts 2, 3 and 4, we look at the cases within the proof of Lemma 2.6. We are interested in the set of components at the end of the game that gives the largest s. In the first case, that is an induced triangle matching, hence $s(n,\#S_{3},P_{4})\leq n-o(n)$. For the second case, that is a disjoint union of $K_{4}$’s, which gives $s(n,\#S_{3},P_{5})\leq 3n-o(n)$ and $s(n,\#S_{4},P_{5})\leq n-o(n)$. In the last case, the graph is a disjoint union of $K_{p}$’s, with $p\leq 5$, and possible a number of specially formed components on $6$ vertices. Hence, for the upper bound we can assume that at the end of the game we have almost a disjoint union of $K_{5}$’s, which gives $s(n,\#S_{3},P_{6})\leq 6n-o(n)$, $s(n,\#S_{4},P_{6})\leq 4n-o(n)$ and $s(n,\#S_{5},P_{6})\leq n-o(n)$.